The pH of a solution is 4.5. What is the concentration of hydroxide ion in the solution.


1
Expert's answer
2021-03-24T01:06:06-0400

According to the definitions:

[OH-] = 10-pOH

pOH = 14 - pH

From here:

pOH = 14 - pH = 14 - 4.5 = 9.5

[OH-] = 10-pOH = 10-9.5 = 3.2 × 10-10 M


Answer: 3.2 × 10-10 M